OpenStudy (luigi0210):

Simple Integral Problem:

OpenStudy (luigi0210):

\[\int\limits \frac{\sin2x}{cosx} dx\]

OpenStudy (psymon):

Use an identity. sin2x = 2sinxcosx

OpenStudy (luigi0210):

So change it to: \[\int\limits \frac{2sinxcosx}{cosx}\]

OpenStudy (luigi0210):

*dx

OpenStudy (abb0t):

simple?! what the heck is this? doesn't look liek any algebra i seen before

OpenStudy (luigi0210):

@abb0t quit playing, you know this.

OpenStudy (psymon):

It's sin is what it is. PURE SIN!

OpenStudy (psymon):

But yes @Luigi0210 , then cosx cancels and you can integrate just fine from there.

OpenStudy (luigi0210):

\[2 \int\limits sinx dx=-2cosx +C\]

OpenStudy (luigi0210):

Thank you @Psymon "_

OpenStudy (luigi0210):

*:)

OpenStudy (abb0t):

I have no idea what's going on, literally.